OpenStudy (anonymous):

Solve the following system of equations. x + y = 15 x – y = -19 A. (2, 17) B. (-2, -17) C. (-2, 17) D. (5, 5)

OpenStudy (anonymous):

need help finding an answer

OpenStudy (anonymous):

Add the two equations together and what do you get?

OpenStudy (anonymous):

x + y = 15 x – y = -19 Luckily, you don't have to use any form of substitution/elimination to solve this. Basically just add them together and divide. :D x + y = 15 x – y = -19 x +x = 2x y -y = 0 15 - 19 = -4 We're left with: 2x = -4 Divide each side by 2. x = -2 <--- We can plug this into one of the original equations to find y. x + y = 15 Plugging in x. -2 + y = 15 Add 2 to each side. y = 17 Solution: (-2, 17) <--- Answer. (:

OpenStudy (anonymous):

You can also plug the solution in to see if it works. (-2, 17) x + y = 15 x – y = -19 -2 + 17 = 15 15 = 15 Yes. x - y = -19 -2 - 17 = -19 -19 = -19 Yes.